OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1).

OpenStudy (anonymous):

always use this forumla \[y-y_{1} =m({x-x_{1}})\] y1 = -1 x1 = 5 m = -2 U get hte equation

OpenStudy (anonymous):

so what would plain y and x be?

OpenStudy (anonymous):

i got the equation but what i got was wrong

OpenStudy (shamim):

u can solve it by another formula

OpenStudy (anonymous):

whats the formula?

OpenStudy (shamim):

y =mx+c

OpenStudy (anonymous):

ok lemme try it out

OpenStudy (shamim):

put in this equation m=-2 x=5 y=-1

OpenStudy (shamim):

c=?

OpenStudy (anonymous):

ok this equation works and c would be either 5,-5,1,or-1

OpenStudy (shamim):

no

OpenStudy (anonymous):

well i looked at the answer choices

OpenStudy (shamim):

y =mx+c -1=(-2)*5+c c=?

OpenStudy (anonymous):

A.y-1=-2(x-5) B.y + 1 = -2(x – 5) C.y -5 = -2(x – 1) D.y – 5 = -2(x + 1)

OpenStudy (shamim):

B

OpenStudy (anonymous):

ok thank you

OpenStudy (shamim):

i think u did not understand it

OpenStudy (anonymous):

how about this question: For $360, a rock-climbing gym offers a yearly membership where members can climb as many days as they want and pay $4 per day for equipment rental. Nonmembers pay $10 per day to use the gym and $6 per day for equipment rental. Find the total number of visits when the cost for a member and nonmember will be the same.

OpenStudy (shamim):

no

OpenStudy (anonymous):

A.174 B.480 C.30 D.60

OpenStudy (shamim):

u r not trying to understand. just collecting result

OpenStudy (anonymous):

well im timed on this test